# Regularity of the global attractor for the 2D incompressible Navier-Stokes equations on channel-like domains

### v1

@@ 2.1. General settings
[…]
@@
We identify $H$ with its dual and consider the dual space $V^{\prime}$ of $V$, so that $V\subseteq H\subseteq V^{\prime}$, with continuous and dense inclusions. For simplicity, we use the notation (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) also for the duality product between $V$ and $V^{\prime}$.
@@
$B(\cdot,\cdot):V\times V\rightarrow D(A^{-s})\text{ is continuous for }0<s\leq\frac{1}{2}.$ (2.16)
@@
in the distribution sense, with $t\mapsto|u(t)|^{2}$ absolutely continuous on [0,)[0$\infty$)[ 0 , ∞ ).
@@
These regularity results follow from enery-type inequalities for $\|u\|$ and $|u_{t}|$, obtained by taking (for the Galerkin approximations) the inner product in $H$ of $u_{t}$ first with (2.10) and then with (2.21),
@@
$\int_{T}^{T+T_{0}}\|u_{t}(t)\|^{2}\;\mathrm{d}t$ $\leq\frac{1}{\tau}C_{2}(T_{0}+\tau,\nu,\lambda_{1},|u(T-\tau)|,\|f\|_{V^{\prime}}),$ $\forall\,T+T_{0}\geq T\geq\tau>0.$
@@ 2.2. More regular domains
for $u\in D(A^{\sigma}),\,v\in V,\,w\in H,$ for any $1/2<\sigma\leq 1$ This means that the bilinear operator is bounded as an operator in
@@ 3. Descriptions of the main results
[…]
@@ 4.1. Regularity of the semigroup
We will show that for positive time $t>0$, the map $S(t)$ is continuous from $H$ into $V$. In fact, it is a Lipschitz map. By subtracting two solution $u(t)$ and $v(t)$, $t\geq 0$, with $u(0)=u_{0}$, $v(0)=v_{0}$, and setting $w(t)=u(t)-v(t)$, we find the equations
@@
[…]
@@
4.2. Compactness of the global attractor in $V$ proof of Theorem 3.1
@@
where $F(u)=f-\nu Au-B(u)$, which by Theorem 3.2 is continuous from ${\mathcal{A}}$ into $H$, when we consider the topology on ${\mathcal{A}}$ to be say $H$.
@@
For $u\in{\mathcal{A}}\subseteq D(A^{\sigma})$, we have that $B(u)$ and $F(u)$ belong to $H$. Moreover $f$ is assumed to be in $H$. Thus, $Au\in H$, which means that ${\mathcal{A}}\subseteq D(A)$. The compactness of ${\mathcal{A}}$ in $D(A)$ also follows from (4.11). Indeed, if $\{u_{n}\}_{n}$ is a sequence in ${\mathcal{A}}$, then, since ${\mathcal{A}}$ is compact in $D(A^{\sigma})$, if follows that $u_{n^{\prime}}\rightarrow u$ in $D(A^{\sigma})$, for some $u\in{\mathcal{A}}$ and for some subsequence $\{u_{n^{\prime}}\}$. By the assumed continuity of $B(\cdot)$ from $D(A^{\sigma})$ into $H$, and by the continuity of $F(\cdot)$ proved in Proposition 3.1, we also have $B(u_{n^{\prime}})+F(u_{n^{\prime}})\rightarrow B(u)+F(u)$ in $H$. Finally, since $f\in H$, it follows from (4.11) that $Au_{n^{\prime}}\rightarrow Au$ in $H$, which proves the desired compactness.
@@ 5. Proofs using energy-type equations
5.1. Compactness of the global attractor in $V$ proof of Theorem 3.1) via the energy method
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}(\lambda_{1}|u|^{2}+\|u\|^{2})+\nu\lambda_{1}\|u\|^{2}+\frac{1}{\nu}|u_{t}|^{2}+\frac{1}{\nu}b(u,u,u_{t})=(f,u)+\frac{1}{\nu}(f,u_{t}).$ (5.1)
@@
[…]
@@
$|u_{n}(t)|^{2}=|u_{0n}|e^{-\nu\lambda_{1}t}+\int_{0}^{t}e^{-\nu\lambda_{1}(t-s)}\left\{K(u_{n}(s),u_{nt}(s))-L(u_{n}(s),u_{nt}(s))\right\}\;\mathrm{d}s,$ (5.4)
@@
$\|u_{n}(t)\|_{1}^{2}=\|u_{n}(t_{n}-T)\|_{1}^{2}e^{-\nu\lambda_{1}T}+\int_{0}^{T}e^{-\nu\lambda_{1}(T-s)}\left\{K(u_{n}(t_{n}-T+s),u_{nt}(t_{n}-T+s))\right.\\ \left.-L(u_{n}(t_{n}-T+s),u_{nt}(t_{n}-T+s))\right\}\;\mathrm{d}s.$ (5.9)
@@
[…]
$\int_{0}^{T}e^{-\nu\lambda_{1}(T-s)}K(S(s)w^{T},\partial_{s}S(s)w^{T})\;\mathrm{d}s\\ =\lim_{n^{\prime}\to\infty}\int_{0}^{T}e^{-\nu\lambda_{1}(T-s)}K(u_{n}(t_{n}-T+s),u_{nt}(t_{n}-T+s))\;\mathrm{d}s.$ (5.13)
@@
$\limsup_{n^{\prime}\to\infty}\|u_{n}(t_{n}-T)\|_{1}^{2}e^{-\nu\lambda_{1}T}\leq\rho_{1}^{2}e^{-\nu\lambda_{1}T}.$ (5.14)
[…]
@@
$|u_{n}^{\prime}(t)|^{2}=|u_{0n}^{\prime}|e^{-2\nu\lambda_{1}t}-2\int_{0}^{t}e^{-2\nu\lambda_{1}(t-s)}\left\{b(u_{n}^{\prime}(s),u_{n}(s),u_{n}^{\prime}(s))+\nu\lambda_{1}[u_{n}^{\prime}(s)]_{1}^{2}\right\}\;\mathrm{d}s,$ (5.23)
@@
$|u_{n}^{\prime}(t)|^{2}=|u_{n}^{\prime}(t_{n}-T)|e^{-2\nu\lambda_{1}T}\\ -2\int_{0}^{T}e^{-2\nu\lambda_{1}(T-s)}\left\{b(u_{n}^{\prime}(t_{n}-T+s),u_{n}(t_{n}-T+s),u_{n}^{\prime}(t_{n}-T+s))\right.\\ \left.+\nu\lambda_{1}[u_{n}^{\prime}(t_{n}-T+s)]_{1}^{2}\right\}\;\mathrm{d}s.$ (5.24)
@@
$|w^{\prime}(0)|^{2}=|w^{\prime}(-T)|e^{-2\nu\lambda_{1}t}\\ -2\int_{0}^{t}e^{-2\nu\lambda_{1}(t-s)}\left\{b(w^{\prime}(s-T),w(s-T),w^{\prime}(s-T))+\nu\lambda_{1}[w^{\prime}(s-T)]_{1}^{2}\right\}\;\mathrm{d}s.$ (5.25)
@@
[…]

### v2

@@ 2.1. General settings
for $u=(u_{1},u_{2}),v=(v_{1},v_{2})\in H_{0}^{1}(\Omega)^{2}.$ The corresponding norms are $|\cdot|=(\cdot,\cdot)^{1/2}$ and $\|\cdot\|=(\!({\cdot,\cdot})\!)^{1/2}$, respectively. Thanks to (2.2) the norm $\|\cdot\|$ is equivalent to the usual one in $H_{0}^{1}(\Omega)^{2}$. Let ${\mathcal{C}}_{\mathrm{c}}^{\infty}(\Omega)$ denote the space of infinitely-differentiable real-valued functions with compact support on $\Omega$. Set
@@
We identify $H$ with its dual and consider the dual space $V^{\prime}$ of $V$, so that $V\subseteq H\subseteq V^{\prime}$, with continuous and dense inclusions. For simplicity, we use the notation $(\cdot,\cdot)$ also for the duality product between $V$ and $V^{\prime}$.
@@
[…]
@@
in the distribution sense, with $t\mapsto|u(t)|^{2}$ absolutely continuous on $0,\infty)$.
@@
These regularity results follow from energy-type inequalities for $\|u\|$ and $|u_{t}|$, obtained by taking (for the Galerkin approximations) the inner product in $H$ of $u_{t}$ first with (2.10) and then with (2.21),
@@
$\int_{T}^{T+T_{0}}\|u_{t}(t)\|^{2}\;\mathrm{d}t$ $\leq\frac{1}{\tau^{2}}C_{2}(T_{0}+\tau,\nu,\lambda_{1},|u(T-\tau)|,\|f\|_{V^{\prime}}),$ $\forall\,T+T_{0}\geq T\geq\tau>0.$
@@ 2.2. More regular domains
for $u\in D(A^{\sigma}),\,v\in V,\,w\in H,$ for any $1/2<\sigma\leq 1.$ This means that the bilinear operator is bounded as an operator in
@@ 3. Descriptions of the main results
Note that from (2.12) the case $s=1/2$ is precisely Theorem 3.1. The extension is for $0<s<1/2$.
@@ 4.1. Regularity of the semigroup
We will show that for positive time $t>0$, the map $S(t)$ is continuous from $H$ into $V$. In fact, it is a Lipschitz map. By subtracting two solution $u(t)$ and $v(t)$, $t\geq 0$, with $u(0)=u_{0}$, $v(0)=v_{0}$, and setting w(t)=u(t)v(t)w(t)=u(t)-v(t), we find the equations
@@
where, for any $T>0$, $\alpha(t)$ is bounded on $[0,T]$ thanks to (2.24), $\beta(t)$ is square-integrable on $[0,T]$ thanks to (2.19) and (2.20), and $\gamma(t)$ is integrable on $[0,T]$ thanks to (2.25) and (2.20). By applying the Gronwall Lemma and using that $y(0)=0$ we find
@@
4.2. Compactness of the global attractor in $V$ - proof of Theorem 3.1
@@
where $F(u)=f-\nu Au-B(u)$, which by Proposition 3.1 is continuous from $\mathcal{A}$} into $H$, when we consider the topology on $\mathcal{A}$} to be say $H$.
@@
For $u\in{\mathcal{A}}\subseteq D(A^{\sigma})$, we have that $B(u)$ and $F(u)$ belong to $H$. Moreover $f$ is assumed to be in $H$. Thus, $Au\in H$, which means that ${\mathcal{A}}\subseteq D(A)$. The compactness of $\mathcal{A}$} in $D(A)$ also follows from (4.11). Indeed, if $\{u_{n}\}_{n}$ is a sequence in $\mathcal{A}$}, then, since $\mathcal{A}$} is compact in $D(A^{\sigma})$, it follows that $u_{n^{\prime}}\rightarrow u$ in $D(A^{\sigma})$, for some $u\in{\mathcal{A}}$ and for some subsequence $\{u_{n^{\prime}}\}$. By the assumed continuity of $B(\cdot)$ from $D(A^{\sigma})$ into $H$, and by the continuity of $F(\cdot)$ proved in Proposition 3.1, we also have $B(u_{n^{\prime}})+F(u_{n^{\prime}})\rightarrow B(u)+F(u)$ in $H$. Finally, since $f\in H$, it follows from (4.11) that $Au_{n^{\prime}}\rightarrow Au$ in $H$, which proves the desired compactness.
@@ 5. Proofs using energy-type equations
5.1. Compactness of the global attractor in $V$ - proof of Theorem 3.1) via the energy method
@@
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}(\lambda_{1}|u|^{2}+\|u\|^{2})+\nu\lambda_{1}\|u\|^{2}+\frac{1}{\nu}|u_{t}|^{2}+\frac{1}{\nu}b(u,u,u_{t})=\lambda_{1}(f,u)+\frac{1}{\nu}(f,u_{t}).$ (5.1)
@@
and set $\|u\|_{1}^{2}=(\!({u,u})\!)_{1}$, $u]_{1}^{2}=[u,u]_{1}$. Clearly, $(\!({\cdot,\cdot})\!)_{1}$, and $\cdot,\cdot]_{1}$ are bilinear and symmetric. Moreover, from the Poincaré inequality (2.3), we see that $\|\cdot\|_{1}$ is a norm in $V$ equivalent to $\|\cdot\|$, while $\cdot\,]_{1}$ is a seminorm in $V$.
@@
$|u_{n}(t)|^{2}=|u_{0n}|^{2}e^{-\nu\lambda_{1}t}+\int_{0}^{t}e^{-\nu\lambda_{1}(t-s)}\left\{K(u_{n}(s),u_{nt}(s))-L(u_{n}(s),u_{nt}(s))\right\}\;\mathrm{d}s,$ (5.4)
@@
$\|u_{n}(t_{n})\|_{1}^{2}=\|u_{n}(t_{n}-T)\|_{1}^{2}e^{-\nu\lambda_{1}T}+\int_{0}^{T}e^{-\nu\lambda_{1}(T-s)}\left\{K(u_{n}(t_{n}-T+s),u_{nt}(t_{n}-T+s))\right.\\ \left.-L(u_{n}(t_{n}-T+s),u_{nt}(t_{n}-T+s))\right\}\;\mathrm{d}s.$ (5.9)
@@
Now we handle the term with $K(\cdot,\cdot)$ in (5.9). From the estimates (2.20) and (2.28) we find that $\{u_{n^{\prime}}(t_{n^{\prime}}-T+\cdot)\}_{n^{\prime}}$ converges weakly to $S(\cdot)w^{T}$ in $W^{1,2}(0,T;V)$, for all $T>0$. Moreover, from the strong convergence (5.7) and the uniform continuity on $[0,T]$ of the semigroup $\{S(t)\}_{t\geq 0}$ in $H$ we see that $\{u_{n^{\prime}}(t_{n^{\prime}}-T+\cdot)\}_{n^{\prime}}$ converges strongly to $S(\cdot)w^{T}$ in $L^{\infty}(0,T;H)$. These are sufficient to yield
$\int_{0}^{T}e^{-\nu\lambda_{1}(T-s)}K(S(s)w^{T},\partial_{s}S(s)w^{T})\;\mathrm{d}s\\ =\lim_{n^{\prime}\to\infty}\int_{0}^{T}e^{-\nu\lambda_{1}(T-s)}K(u_{n^{\prime}}(t_{n^{\prime}}-T+s),u_{n^{\prime}t}(t_{n^{\prime}}-T+s))\;\mathrm{d}s.$ (5.13)
@@
$\limsup_{n^{\prime}\to\infty}\|u_{n^{\prime}}(t_{n^{\prime}}-T)\|_{1}^{2}e^{-\nu\lambda_{1}T}\leq\rho_{1}^{2}e^{-\nu\lambda_{1}T}.$ (5.14)
Taking the $\limsup$ as $n^{\prime}$ goes to infinity in (5.9) and using the previous relations,
@@
$|u_{n}^{\prime}(t)|^{2}=|u_{0n}^{\prime}|^{2}e^{-2\nu\lambda_{1}t}-2\int_{0}^{t}e^{-2\nu\lambda_{1}(t-s)}\left\{b(u_{n}^{\prime}(s),u_{n}(s),u_{n}^{\prime}(s))+\nu\lambda_{1}[u_{n}^{\prime}(s)]_{1}^{2}\right\}\;\mathrm{d}s,$ (5.23)
@@
$|u_{n}^{\prime}(t)|^{2}=|u_{n}^{\prime}(t_{n}-T)|^{2}e^{-2\nu\lambda_{1}T}\\ -2\int_{0}^{T}e^{-2\nu\lambda_{1}(T-s)}\left\{b(u_{n}^{\prime}(t_{n}-T+s),u_{n}(t_{n}-T+s),u_{n}^{\prime}(t_{n}-T+s))\right.\\ \left.+\nu\lambda_{1}[u_{n}^{\prime}(t_{n}-T+s)]_{1}^{2}\right\}\;\mathrm{d}s.$ (5.24)
@@
$|w^{\prime}(0)|^{2}=|w^{\prime}(-T)|^{2}e^{-2\nu\lambda_{1}t}\\ -2\int_{0}^{t}e^{-2\nu\lambda_{1}(t-s)}\left\{b(w^{\prime}(s-T),w(s-T),w^{\prime}(s-T))+\nu\lambda_{1}[w^{\prime}(s-T)]_{1}^{2}\right\}\;\mathrm{d}s.$ (5.25)
@@
Taking the $\limsup$ as $n$ goes to infinity in (5.24) and using the previous relations and the estimate (5.18), we obtain